per donut is a daily cost. The cost of producing the first donut of the day is $5, the second donut of the day is $6 and so on. This holds for everyday Peter decides to produce.
Peter is an architect who currently makes $15 a day in wages. He is considering leaving his job and opening a donut shop. The cost of producing donuts is $5 for the first donut and increases by $I for each subsequent donut produced (i.e. the first donut costs $5 to produce, the second donut costs $6 to produce, the third donut costs $7 to produce, and so on). In addition, Peter must pay a daily cost of $5 to operate the shop, regardless of the amount of donuts he produces. Peter would be able to sell each donut for $10. If Peter only cares about his income, what should he do? Explain.
Answer #1

Answer

Lets calculate how much Profit(Accounting) he can earn If he open a donut shop. We can see that Cost of producing donut increases by $1 as it produces 1 subsequent donut. Cost of producing 7th donut = MC for producing 7th donut = 11 and MC for 6th donut = 10. Also Price = MR for 6th unit = 10.

According to profit maximizing condition a firm produces that quantity at which MR = MC

Hence they will produce 6 units Because at 6th Unit MR = MC = 10.

Hence total Accounting Profit earned = 10*6 - 45 - 5 = 10

where 45 = Variable Cost = 5 + 6 + 7 + 8 + 9 + 10

Fixed Cost = 5(per day rent)

If he works instead of opening the job.he will earn $15 which is greater than prodfit earned if he opens the job( = $15). Hence as he cares only about profit he should work isntead of opening the jod.

Hence, Peter should works as an architect earn $15